# Reinforcement learning control of quantum error correction

### v2

@@ Abstract
We have thus far demonstrated that the mean $\mu(t)$ of the Gaussian policy distribution learns to track the optimal policy over time in the presence of drift. As a result, the learned policy $\mu(t)$ has superior performance compared to the fixed policy, with lower DR and LER. However, since the learning process requires exploration of the parameter space, the algorithm inevitably samples policy candidates whose performance is worse than that of the $\mu(t)$ policy. This “exploration noise” is irrelevant in our experimental setting where RL steering relies on repetitive executions of a short logical algorithm and the quantum state is independently re-prepared in every shot. However, in the future this steering must be done in real time during the single-shot execution of a long logical algorithm. In that case, the exploration noise, although necessary for learning, is going to be detrimental to the logical algorithm’s performance. Generally, balancing the exploration of parameter space and exploitation of the learned policy $\mu(t)$ is a central challenge in many applications of RL [sutton2018reinforcement]. In our case, the favorable resolution of the exploration-exploitation tradeoff would mean that aggregate performance of all sampled policy candidates, most of which are worse than $\mu(t)$, is still better than the performance without steering.
@@
To study this tradeoff, we conducted numerical simulations of real-time steering of the distance-3 surface code subject to sinusoidal parameter drift at different frequencies, see Fig. 5(a) and Supp. Mat. V A. We count the total number of error detection events generated in a $1.8\times 10^{9}$-cycle window of QEC, and normalize it so that level 1 corresponds to performance of the optimal policy (known in the simulation), and level 0 corresponds to the performance of a fixed policy (optimal at $t=0$). We control the exploration-exploitation tradeoff by changing the amount of entropy regularization [haarnoja2018soft] in our RL algorithm. Our findings indicate that there is a critical drift frequency, about $1/150$ epochs, below which the system becomes real-time steerable: the performance with exploration noise is better than the performance of a fixed policy. This critical frequency is consistent with the response time of the learning algorithm observed in experiment, see Fig. 4(b). When drift is too fast, the real-time steering is not able to keep up – such drift must be mitigated at the hardware level. In contrast, at low drift frequencies, the exploration and exploitation can be successfully balanced to closely approach the performance of the optimal policy at all times.
@@
Thus, our simulations establish that RL is able to effectively utilize the information concealed in the error detection events to calibrate the system while continuing the logical computation. This ability offers a significant advantage over approaches based on the synthesis of traditional calibration and code deformation [fang2024caliscalpel], since RL steering does not introduce any resource overhead.
@@
Figure 5: Real-time steering and scaling simulations. (a) Normalized improvement (color) of detection rate in the real-time steering simulation of the distance-3 surface code subject to sinusoidal drift at different frequencies. Level 11 indicates the performance of optimal policy. Isoline at level 0 demarcates the boundary beyond which real-time steering results in better performance than a fixed policy, approaching the performance of optimal policy in the regime of slow drift. (b) Simulation of scalability of RL control of large surface codes. The algorithm learns the parameters of single-qubit and CZ gates, with $30$ control parameters per gate, amounting to almost $40,000$ control parameters in total for the distance-15 code. During the learning process, the logical error rate reduces over time (color) until it reaches the floor (red bars) set by the irreducible physical error rates and characterized by the error suppression factor $\Lambda^{*}$. (c) Point-estimates of $\Lambda$ at every code distance and learning epoch from (b) confirm that the speed, $\partial_{t}\Lambda/\Lambda^{*}\times 10^{2}$, at which the error suppression factor approaches the local optimum, is proportional to the distance from the optimum, $1-\Lambda/\Lambda^{*}$. The convergence rate $\gamma$ is independent of the system size but depends on the number of control parameters per gate, with three beams corresponding to 1, 10, and 30 parameters, and the linear fits (red) indicating the convergence rates.
@@
Finally, we demonstrate the scalability of RL control with simulations extending to a distance-15 surface code with approximately $40,000$ control parameters, see Fig. 5(b). We extract point-estimates of the “instantaneous” error suppression factor $\Lambda$ at each code distance and learning epoch, verifying in Fig. 5(c) that its convergence rate $\partial_{t}\Lambda/\Lambda^{*}$ is proportional to the distance from the optimal $\Lambda^{*}$, set in our simulation by the irreducible control-independent physical error rates. This empirical result leads to the exponential convergence law, $1-\Lambda/\Lambda^{*}\propto e^{-\gamma\,t}$, characteristic of gradient-descent based algorithms in the vicinity of local optimum [boyd2004convex], with the convergence rate $\gamma$ that depends on properties such as the number of gates per detecting region and the number of control parameters per gate. Crucially, this rate is independent of the system size, a direct consequence of our algorithm’s ability to effectively utilize the sparsity of the factor graph representation of the control problem.
@@
Multiple possible extensions of our RL control framework leave ample room to further enhance its capabilities. For example, instead of relying on a simple Gaussian policy distribution, it could be advantageous to employ deep neural networks [mnih2015human] and condition the policy distribution on the observations from the environment, e.g. on some suitable statistics of the error detection events. Learning the system model in addition to the policy [hafner2019dream] can lead to better sample efficiency and further improve the quality of solutions by discovering unexpected correlations in the data. It is conceivable that in the future, with sufficient enhancement of our learning framework, a quantum processor could be calibrated for QEC ab initio fully by RL, with no reliance on the traditional calibration paradigm or human experts.
@@
In conclusion, by empowering the quantum computer to learn directly from its errors, we unlock a scalable pathway to optimize performance in real time, replacing disruptive calibration routines with uninterrupted computation. Our work suggests that the path to fault tolerance will be built not only on better hardware, but on more intelligent control.
@@
V.S. conceived and led the project, with significant contribution from A.M; V.S. developed the RL control framework; A.M. calibrated and expert-tuned the processor; V.S. and A.M. designed and conducted the experiments and wrote the manuscript with input from all the authors; M.B. and M.N. contributed software improvements; A.E. performed device simulations; P.K. supervised the project. All authors contributed to building and maintaining the hardware, software, cryogenics, and electronics infrastructure.
@@
Google Quantum AI and Collaborators
@@
Volodymyr Sivak1, Alexis Morvan1, Michael Broughton1, Matthew Neeley1, Alec Eickbusch1, Dmitry Abanin1, Amira Abbas1, Rajeev Acharya1, Laleh Aghababaie Beni1, Georg Aigeldinger1, Ross Alcaraz1, Sayra Alcaraz1, Trond I. Andersen1, Markus Ansmann1, Frank Arute1, Kunal Arya1, Walt Askew1, Nikita Astrakhantsev1, Juan Atalaya1, Brian Ballard1, Joseph C. Bardin1, 2, Hector Bates1, Andreas Bengtsson1, Majid Bigdeli Karimi1, Alexander Bilmes1, Simon Bilodeau1, Felix Borjans1, Alexandre Bourassa1, Jenna Bovaird1, Dylan Bowers1, Leon Brill1, Peter Brooks1, David A. Browne1, Brett Buchea1, Bob B. Buckley1, Tim Burger1, Brian Burkett1, Nicholas Bushnell1, Jamal Busnaina1, Anthony Cabrera1, Juan Campero1, Hung-Shen Chang1, Silas Chen1, Ben Chiaro1, Liang-Ying Chih1, Agnetta Y. Cleland1, Bryan Cochrane1, Matt Cockrell1, Josh Cogan1, Roberto Collins1, Paul Conner1, Harold Cook1, Rodrigo G. Cortiñas1, William Courtney1, Alexander L. Crook1, Ben Curtin1, Martin Damyanov1, Sayan Das1, Dripto M. Debroy1, Sean Demura1, Paul Donohoe1, Ilya Drozdov1, Andrew Dunsworth1, Valerie Ehimhen1, Aviv Moshe Elbag1, Lior Ella1, Mahmoud Elzouka1, David Enriquez1, Catherine Erickson1, Vinicius S. Ferreira1, Marcos Flores1, Leslie Flores Burgos1, Ebrahim Forati1, Jeremiah Ford1, Austin G. Fowler1, Brooks Foxen1, Masaya Fukami1, Alan Wing Lun Fung1, Lenny Fuste1, Suhas Ganjam1, Gonzalo Garcia1, Christopher Garrick1, Robert Gasca1, Helge Gehring1, Robert Geiger1, Élie Genois1, William Giang1, Dar Gilboa1, James E. Goeders1, Edward C. Gonzales1, Raja Gosula1, Stijn J. de Graaf1, Alejandro Grajales Dau1, Dietrich Graumann1, Joel Grebel1, Alex Greene1, Jonathan A. Gross1, Jose Guerrero1, Loïck Le Guevel1, Tan Ha1, Steve Habegger1, Tanner Hadick1, Ali Hadjikhani1, Michael C. Hamilton1, Matthew P. Harrigan1, Sean D. Harrington1, Jeanne Hartshorn1, Stephen Heslin1, Paula Heu1, Oscar Higgott1, Reno Hiltermann1, Hsin-Yuan Huang1, Mike Hucka1, Christopher Hudspeth1, Ashley Huff1, William J. Huggins1, Evan Jeffrey1, Shaun Jevons1, Zhang Jiang1, Xiaoxuan Jin1, Chaitali Joshi1, Pavol Juhas1, Andreas Kabel1, Dvir Kafri1, Hui Kang1, Kiseo Kang1, Amir H. Karamlou1, Ryan Kaufman1, Kostyantyn Kechedzhi1, Tanuj Khattar1, Mostafa Khezri1, Seon Kim1, Can M. Knaut1, Bryce Kobrin1, Fedor Kostritsa1, John Mark Kreikebaum1, Ryuho Kudo1, Ben Kueffler1, Arun Kumar1, Vladislav D. Kurilovich1, Vitali Kutsko1, Nathan Lacroix1, David Landhuis1, Tiano Lange-Dei1, Brandon W. Langley1, Pavel Laptev1, Kim-Ming Lau1, Justin Ledford1, Joy Lee1, Kenny Lee1, Brian J. Lester1, Wendy Leung1, Lily Li1, Wing Yan Li1, Ming Li1, Alexander T. Lill1, William P. Livingston1, Matthew T. Lloyd1, Aditya Locharla1, Laura De Lorenzo1, Daniel Lundahl1, Aaron Lunt1, Sid Madhuk1, Aniket Maiti1, Ashley Maloney1, Salvatore Mandrà1, Leigh S. Martin1, Orion Martin1, Eric Mascot1, Paul Masih Das1, Dmitri Maslov1, Melvin Mathews1, Cameron Maxfield1, Jarrod R. McClean1, Matt McEwen1, Seneca Meeks1, Kevin C. Miao1, Zlatko K. Minev1, Reza Molavi1, Sebastian Molina1, Shirin Montazeri1, Charles Neill1, Michael Newman1, Anthony Nguyen1, Murray Nguyen1, Chia-Hung Ni1, Murphy Yuezhen Niu1, Logan Oas1, Raymond Orosco1, Kristoffer Ottosson1, Alice Pagano1, Agustin Di Paolo1, Sherman Peek1, David Peterson1, Alex Pizzuto1, Elias Portoles1, Rebecca Potter1, Orion Pritchard1, Michael Qian1, Chris Quintana1, Arpit Ranadive1, Matthew J. Reagor1, Rachel Resnick1, David M. Rhodes1, Daniel Riley1, Gabrielle Roberts1, Roberto Rodriguez1, Emma Ropes1, Lucia B. De Rose1, Eliott Rosenberg1, Emma Rosenfeld1, Dario Rosenstock1, Elizabeth Rossi1, Pedram Roushan1, David A. Rower1, Robert Salazar1, Kannan Sankaragomathi1, Murat Can Sarihan1, Kevin J. Satzinger1, Max Schaefer1, 3, Sebastian Schroeder1, Henry F. Schurkus1, Aria Shahingohar1, Michael J. Shearn1, Aaron Shorter1, Noah Shutty1, Vladimir Shvarts1, Spencer Small1, W. Clarke Smith1, David A. Sobel1, Barrett Spells1, Sofia Springer1, George Sterling1, Jordan Suchard1, Aaron Szasz1, Alexander Sztein1, Madeline Taylor1, Jothi Priyanka Thiruraman1, Douglas Thor1, Dogan Timucin1, Eifu Tomita1, Alfredo Torres1, M. Mert Torunbalci1, Hao Tran1, Abeer Vaishnav1, Justin Vargas1, Sergey Vdovichev1, Guifre Vidal1, Catherine Vollgraff Heidweiller1, Meghan Voorhees1, Steven Waltman1, Jonathan Waltz1, Shannon X. Wang1, Brayden Ware1, James D. Watson1, Yonghua Wei1, Travis Weidel1, Theodore White1, Kristi Wong1, Bryan W. K. Woo1, Christopher J. Wood1, Maddy Woodson1, Cheng Xing1, Z. Jamie Yao1, Ping Yeh1, Bicheng Ying1, Juhwan Yoo1, Noureldin Yosri1, Elliot Young1, Grayson Young1, Adam Zalcman1, Ran Zhang1, Yaxing Zhang1, Ningfeng Zhu1, Nicholas Zobrist1, Zhenjie Zou1, Ryan Babbush1, Dave Bacon1, Sergio Boixo1, Yu Chen1, Zijun Chen1, Michel Devoret1, 3, Monica Hansen1, Jeremy Hilton1, Cody Jones1, Julian Kelly1, Alexander N. Korotkov1, Erik Lucero1, Anthony Megrant1, Hartmut Neven1, William D. Oliver1, Ganesh Ramachandran1, Vadim Smelyanskiy1, Paul V. Klimov1
@@
$^{\textrm{\scriptsize 2}}$ Department of Electrical and Computer Engineering, University of Massachusetts, Amherst, MA
$^{\textrm{\scriptsize 3}}$ Department of Physics, University of California, Santa Barbara, CA

### v3

@@ Abstract
We have thus far demonstrated that the mean $\mu(t)$ of the Gaussian policy distribution learns to track the optimal policy over time in the presence of drift. As a result, the learned policy $\mu(t)$ has superior performance compared to the fixed policy, with lower EDR and LER. However, since the learning process requires exploration of the parameter space, the algorithm inevitably samples policy candidates whose performance is worse than that of the $\mu(t)$ policy. This “exploration noise” is irrelevant in our experimental setting where RL steering relies on repetitive executions of a short logical algorithm and the quantum state is independently re-prepared in every shot. However, in the future this steering must be done in real time during the single-shot execution of a long logical algorithm. In that case, the exploration noise, although necessary for learning, is going to be detrimental to the logical algorithm’s performance. Generally, balancing the exploration of parameter space and exploitation of the learned policy $\mu(t)$ is a central challenge in many applications of RL [51]. In our case, the favorable resolution of the exploration-exploitation tradeoff would mean that aggregate performance of all sampled policy candidates, most of which are worse than $\mu(t)$, is still better than the performance without RL steering.
@@
To study this tradeoff, we conducted numerical simulations of real-time steering of the distance-3 surface code subject to sinusoidal parameter drift at different frequencies, see Fig. 5(a) and Supp. Mat. VI A. We count the total number of error detection events generated in a $1.8\times 10^{9}$-cycle window of QEC, and normalize it so that level 1 corresponds to performance of the optimal policy (known in the simulation), and level 0 corresponds to the performance of a fixed policy (optimal at $t=0$). We control the exploration-exploitation tradeoff by changing the amount of entropy regularization [18] in our RL algorithm. Our findings indicate that there is a critical drift frequency, about $1/150$ epochs, below which the system becomes real-time steerable: the performance with exploration noise is better than the performance of a fixed policy. This critical frequency is consistent with the response time of the learning algorithm observed in experiment, see Fig. 4(b). When drift is too fast, the real-time steering is not able to keep up – such drift must be mitigated at the hardware level. In contrast, at low drift frequencies, the exploration and exploitation can be successfully balanced to closely approach the performance of the optimal policy at all times.
@@
Thus, our simulations establish that RL is able to effectively utilize the information concealed in the error detection events to calibrate the system while continuing the logical computation. This ability offers a significant advantage over approaches based on the synthesis of traditional calibration and code deformation [10], since RL steering does not introduce any resource overhead.
@@
Figure 5: Real-time steering and scaling simulations. (a) Normalized improvement (color) of error detection rate in the real-time steering simulation of the distance-3 surface code subject to sinusoidal drift at different frequencies. Level 11 indicates the performance of optimal policy. Isoline at level 0 demarcates the boundary beyond which real-time steering results in better performance than a fixed policy, approaching the performance of optimal policy in the regime of slow drift. (b) Simulation of scalability of RL control of large surface codes. The algorithm learns the parameters of single-qubit and CZ gates, with $30$ control parameters per gate, amounting to almost $40,000$ control parameters in total for the distance-15 code. During the learning process, the logical error rate reduces over time (color) until it reaches the floor (red bars) set by the irreducible physical error rates and characterized by the optimal error suppression factor $\Lambda^{*}$. (c) Point-estimates of $\Lambda$ at every code distance and learning epoch from (b) confirm that the speed, $\partial_{t}\Lambda/\Lambda^{*}\times 10^{2}$, at which the error suppression factor approaches the local optimum, is proportional to the distance from the optimum, $1-\Lambda/\Lambda^{*}$. The convergence rate $\gamma$ (see main text) is independent of the system size but depends on the number of control parameters per gate, with three beams corresponding to 1, 10, and 30 parameters, and the linear fits (red) indicating the convergence rates.
@@
Finally, we demonstrate the scalability of RL control with simulations extending to a distance-15 surface code with approximately $40,000$ control parameters, see Fig. 5(b). We extract point-estimates of the instantaneous error suppression factor $\Lambda$ at each code distance and learning epoch, verifying in Fig. 5(c) that its convergence rate $\partial_{t}\Lambda/\Lambda^{*}$ is proportional to the distance from the optimal $\Lambda^{*}$, set in our simulation by the irreducible control-independent physical error rates. This empirical result leads to the exponential convergence law, $1-\Lambda/\Lambda^{*}\propto e^{-\gamma\,t}$, characteristic of gradient-descent based algorithms in the vicinity of local optimum [5], with the convergence rate $\gamma$ that depends on properties such as the number of gates per detecting region and the number of control parameters per gate. Crucially, this rate is independent of the system size, a direct consequence of our algorithm’s ability to effectively utilize the sparsity of the factor graph representation of the control problem.
@@
Multiple possible extensions of our RL control framework leave ample room to further enhance its capabilities. For example, instead of relying on a simple Gaussian policy distribution, it could be advantageous to employ deep neural networks [34] and condition the policy distribution on the observations from the environment, e.g. on some suitable statistics of the error detection events. Learning the system model in addition to the policy [19] can lead to better sample efficiency and further improve the quality of solutions by discovering unexpected correlations in the data. It is conceivable that in the future, with sufficient enhancement of our learning framework, a quantum processor could be calibrated for QEC ab initio fully by RL, with no reliance on the traditional calibration paradigm or human experts.
@@
In conclusion, by empowering the quantum computer to learn from its errors, we unlock a scalable pathway to optimize performance in real time, replacing disruptive calibration routines with uninterrupted computation. Our work suggests that the path to fault tolerance will be built not only on better hardware, but on more intelligent control.
@@
V.S. conceived and led the project, with significant contribution from A.M; V.S. developed the RL control framework; A.M. and R.G.C. calibrated and expert-tuned the processor; V.S. and A.M. designed and conducted the experiments and wrote the manuscript with input from all authors; M.B. and M.N. contributed software improvements; N.S., L.A.B. and DeepMind team contributed high-accuracy decoding; A.E. performed device simulations; P.K. supervised the project. All authors contributed to building and maintaining the hardware, software, cryogenics, and electronics infrastructure.
@@
Methods
Objective function
Using LER as an optimization objective faces fundamental obstacles to scalability, particularly in multi-qubit stabilizer codes such as the surface code, for the following reasons: (i) The LER is suppressed with the code distance $d$ as $\varepsilon_{\rm L}\propto\Lambda^{-d/2}$, where the error suppression factor $\Lambda>2$ has already been demonstrated in experiments with superconducting circuits [17] and neutral atoms [4]. Thus, accurately resolving the LER would require an exponentially large number of QEC cycles. (ii) The optimization involves a vast number of parameters – already more than two thousand in our $d=7$ experiment, and scaling as $O(d^{2})$ – which renders global optimization from a single scalar metric impractical. (iii) The LER is unsuitable for real-time calibration and steering during a quantum computation, as the logical state is generally unknown. While direct LER optimization has proven effective in specific contexts, such as for small bosonic codes [48], these scaling challenges necessitate a different approach for large multi-qubit codes.
The surrogate objective $C$, defined in the main text, immediately alleviates the limitation (iii). It also alleviates (i), since resolving $C$ to a constant relative accuracy requires a number of QEC cycles that scales as $O(\varepsilon^{-1})$ and is independent of the code distance. To alleviate (ii) and enable efficient high-dimensional optimization, we harness the sparse structure of the surrogate objective. We rely on the locality of detectors in the QEC circuit to narrow down the dependence of each term in $C$ to only a small subset of control parameters – those related to the gates within the respective detecting region. These relations can be conveniently represented using a factor graph, a bipartite graph in which the factor nodes correspond to detectors, and the variable nodes correspond to the system control parameters, see Fig. 2(d) and Supp. Mat. II for more information.
Learning algorithm
Directly optimizing the surrogate objective $C$ is intractable for standard off-the-shelf optimizers. Although some algorithms [52] are compatible with large-scale nonlinear optimization problems represented as sparse factor graphs, in practice additional considerations play an important role. Specifically, our objective function $C$ is highly stochastic, as it is obtained by averaging the binary detector values from a finite (and preferably small) number of QEC cycles. Furthermore, $C$ is non-stationary due to inherent system drift. Additionally, evaluating $C$ is expensive, since it requires updating the classical controller and acquiring data from the QEC detectors. Finally, in the future, $C$ must be optimized in real time during the quantum computation, favoring smooth and gradual improvements over disruptive large parameter steps that might lead to catastrophic consequences for the logical algorithm.
Given these properties, we adopted a multi-objective policy-gradient reinforcement learning approach, in which signals from QEC detectors comprising $C$ serve as multiple objectives optimized simultaneously. Our learning algorithm is built on parameter-exploring policy gradients [42], an approach where an entire control policy is sampled in one piece, which provides a convenient interface with the classical controller. For robustness in the stochastic and non-stationary setting, we integrate several key techniques: proximal policy optimization [41] for stability, entropy regularization [18] to maintain exploration, and a replay buffer [34] for improved sample efficiency. Crucially, the algorithm harnesses the sparse structure of the factor graph representation of $C$ for efficient variance reduction of the Monte Carlo gradient estimator [35] via gradient masking. The latter technique is adopted from Appendix G of Ref. [47], where it was derived from the point of view of multi-agent learning.
Google Quantum AI and Google DeepMind
@@ Methods
Volodymyr Sivak1, Alexis Morvan1, Michael Broughton1, Rodrigo G. Cortiñas1, Johannes Bausch2, Andrew W. Senior2, Matthew Neeley1, Alec Eickbusch1, Noah Shutty1, Laleh Aghababaie Beni1, James S. Spencer2, Francisco J. H Heras2, Thomas Edlich2, Dmitry Abanin1, Amira Abbas1, Rajeev Acharya1, Georg Aigeldinger1, Ross Alcaraz1, Sayra Alcaraz1, Trond I. Andersen1, Markus Ansmann1, Frank Arute1, Kunal Arya1, Walt Askew1, Nikita Astrakhantsev1, Juan Atalaya1, Brian Ballard1, Joseph C. Bardin1, 3, Hector Bates1, Andreas Bengtsson1, Majid Bigdeli Karimi1, Alexander Bilmes1, Simon Bilodeau1, Felix Borjans1, Alexandre Bourassa1, Jenna Bovaird1, Dylan Bowers1, Leon Brill1, Peter Brooks1, David A. Browne1, Brett Buchea1, Bob B. Buckley1, Tim Burger1, Brian Burkett1, Nicholas Bushnell1, Jamal Busnaina1, Anthony Cabrera1, Juan Campero1, Hung-Shen Chang1, Silas Chen1, Ben Chiaro1, Liang-Ying Chih1, Agnetta Y. Cleland1, Bryan Cochrane1, Matt Cockrell1, Josh Cogan1, Roberto Collins1, Paul Conner1, Harold Cook1, William Courtney1, Alexander L. Crook1, Ben Curtin1, Martin Damyanov1, Sayan Das1, Dripto M. Debroy1, Sean Demura1, Paul Donohoe1, Ilya Drozdov1, Andrew Dunsworth1, Valerie Ehimhen1, Aviv Moshe Elbag1, Lior Ella1, Mahmoud Elzouka1, David Enriquez1, Catherine Erickson1, Vinicius S. Ferreira1, Marcos Flores1, Leslie Flores Burgos1, Ebrahim Forati1, Jeremiah Ford1, Austin G. Fowler1, Brooks Foxen1, Masaya Fukami1, Alan Wing Lun Fung1, Lenny Fuste1, Suhas Ganjam1, Gonzalo Garcia1, Christopher Garrick1, Robert Gasca1, Helge Gehring1, Robert Geiger1, Élie Genois1, William Giang1, Dar Gilboa1, James E. Goeders1, Edward C. Gonzales1, Raja Gosula1, Stijn J. de Graaf1, Alejandro Grajales Dau1, Dietrich Graumann1, Joel Grebel1, Alex Greene1, Jonathan A. Gross1, Jose Guerrero1, Loïck Le Guevel1, Tan Ha1, Steve Habegger1, Tanner Hadick1, Ali Hadjikhani1, Michael C. Hamilton1, Matthew P. Harrigan1, Sean D. Harrington1, Jeanne Hartshorn1, Stephen Heslin1, Paula Heu1, Oscar Higgott1, Reno Hiltermann1, Hsin-Yuan Huang1, Mike Hucka1, Christopher Hudspeth1, Ashley Huff1, William J. Huggins1, Evan Jeffrey1, Shaun Jevons1, Zhang Jiang1, Xiaoxuan Jin1, Chaitali Joshi1, Pavol Juhas1, Andreas Kabel1, Dvir Kafri1, Hui Kang1, Kiseo Kang1, Amir H. Karamlou1, Ryan Kaufman1, Kostyantyn Kechedzhi1, Tanuj Khattar1, Mostafa Khezri1, Seon Kim1, Can M. Knaut1, Bryce Kobrin1, Fedor Kostritsa1, John Mark Kreikebaum1, Ryuho Kudo1, Ben Kueffler1, Arun Kumar1, Vladislav D. Kurilovich1, Vitali Kutsko1, Nathan Lacroix1, David Landhuis1, Tiano Lange-Dei1, Brandon W. Langley1, Pavel Laptev1, Kim-Ming Lau1, Justin Ledford1, Joy Lee1, Kenny Lee1, Brian J. Lester1, Wendy Leung1, Lily Li1, Wing Yan Li1, Ming Li1, Alexander T. Lill1, William P. Livingston1, Matthew T. Lloyd1, Aditya Locharla1, Laura De Lorenzo1, Daniel Lundahl1, Aaron Lunt1, Sid Madhuk1, Aniket Maiti1, Ashley Maloney1, Salvatore Mandrà1, Leigh S. Martin1, Orion Martin1, Eric Mascot1, Paul Masih Das1, Dmitri Maslov1, Melvin Mathews1, Cameron Maxfield1, Jarrod R. McClean1, Matt McEwen1, Seneca Meeks1, Kevin C. Miao1, Zlatko K. Minev1, Reza Molavi1, Sebastian Molina1, Shirin Montazeri1, Charles Neill1, Michael Newman1, Anthony Nguyen1, Murray Nguyen1, Chia-Hung Ni1, Murphy Yuezhen Niu1, Logan Oas1, Raymond Orosco1, Kristoffer Ottosson1, Alice Pagano1, Agustin Di Paolo1, Sherman Peek1, David Peterson1, Alex Pizzuto1, Elias Portoles1, Rebecca Potter1, Orion Pritchard1, Michael Qian1, Chris Quintana1, Arpit Ranadive1, Matthew J. Reagor1, Rachel Resnick1, David M. Rhodes1, Daniel Riley1, Gabrielle Roberts1, Roberto Rodriguez1, Emma Ropes1, Lucia B. De Rose1, Eliott Rosenberg1, Emma Rosenfeld1, Dario Rosenstock1, Elizabeth Rossi1, Pedram Roushan1, David A. Rower1, Robert Salazar1, Kannan Sankaragomathi1, Murat Can Sarihan1, Kevin J. Satzinger1, Max Schaefer1, 4, Sebastian Schroeder1, Henry F. Schurkus1, Aria Shahingohar1, Michael J. Shearn1, Aaron Shorter1, Vladimir Shvarts1, Spencer Small1, W. Clarke Smith1, David A. Sobel1, Barrett Spells1, Sofia Springer1, George Sterling1, Jordan Suchard1, Aaron Szasz1, Alexander Sztein1, Madeline Taylor1, Jothi Priyanka Thiruraman1, Douglas Thor1, Dogan Timucin1, Eifu Tomita1, Alfredo Torres1, M. Mert Torunbalci1, Hao Tran1, Abeer Vaishnav1, Justin Vargas1, Sergey Vdovichev1, Guifre Vidal1, Catherine Vollgraff Heidweiller1, Meghan Voorhees1, Steven Waltman1, Jonathan Waltz1, Shannon X. Wang1, Brayden Ware1, James D. Watson1, Yonghua Wei1, Travis Weidel1, Theodore White1, Kristi Wong1, Bryan W. K. Woo1, Christopher J. Wood1, Maddy Woodson1, Cheng Xing1, Z. Jamie Yao1, Ping Yeh1, Bicheng Ying1, Juhwan Yoo1, Noureldin Yosri1, Elliot Young1, Grayson Young1, Adam Zalcman1, Ran Zhang1, Yaxing Zhang1, Ningfeng Zhu1, Nicholas Zobrist1, Zhenjie Zou1, Ryan Babbush1, Dave Bacon1, Sergio Boixo1, Yu Chen1, Zijun Chen1, Michel Devoret1, 4, Monica Hansen1, Jeremy Hilton1, Cody Jones1, Julian Kelly1, Alexander N. Korotkov1, Erik Lucero1, Anthony Megrant1, Hartmut Neven1, William D. Oliver1, Ganesh Ramachandran1, Vadim Smelyanskiy1, Paul V. Klimov1
@@
$^{\textrm{\scriptsize 2}}$ Google DeepMind
$^{\textrm{\scriptsize 3}}$ Department of Electrical and Computer Engineering, University of Massachusetts, Amherst, MA
$^{\textrm{\scriptsize 4}}$ Department of Physics, University of California, Santa Barbara, CA